# Collaborative Document Editing with Multiple Users and AI Agents

### v1

@@ 6.5.1. Agent profiles
Tasks were most frequently concerned with revising a document (38), followed by drafting (20) and outlining (5). Users’ primary intents in task descriptions were efficiency (32), improving the text / readability (22), seeking inspiration (18), and learning (15). The mean text length of task descriptions was 12.46 words (SD=14.04, range: 1-102).
@@
Tasks in comments most frequently concerned drafting (36) and revising (34), and less frequently outlining (4). Users’ primary intents in comment delegations were inspiration (35), improving the text / readability (24), efficiency (21), and learning (17). The mean text length was 12.23 words (SD=11.75, range: 0-66). Note that we removed the mentioning of agents before computing length (e.g. “@aiRephraser”).
@@ 7. Discussion
We discuss the findings with respect to the two research questions and further aspects. Figure 6 provides a high-level overview.
@@ 7.1.1. Agent profiles as territory: Each agent belongs to its creator
As a key insight, although our design allowed anyone to edit any agent profile (Section 4), participants almost never did so (Section 6.3.1). Instead, they treated agents as belonging to their creator, which is a form of territorial functioning (Larsen-Ledet and Korsgaard, 2019). Rationales included individual familiarity, alignment with personal writing style, and emerging norms for co-creating agents (Section 6.4.1).
This behaviour echoes prior work on collaboration avoidance, such as role dynamics and accountability and credit of contribution described by Wang et al. (2017). Related concerns in our study included the perceived roles of “creator” and “user” for each agent, or preferring to create a new agent rather than modify one contributed by someone else. Related, Larsen-Ledet and Korsgaard (2019) describe linkage with place, including local expertise, as a dimension of writing territoriality. In this light, our observed behaviour may suggest that the creator of an agent is respected as its local expert by the team.
Finally, Gero et al. (2023) identified availability, trust, and individuality as aspects that shape a writer’s perception of a support actor (human or AI). Through this lens, our findings imply that our participants shared these values and transferred them to the new setting of a shared AI support actor: While agents were available for anyone to modify, participants included individual familiarity, representativity, and trust (e.g. in the sense of “I would not like [others to edit it] if I’m not seeing this”, $P_{2,7}$) in their rationales for “owning” an agent and respecting such ownership of others.
These insights invite future work to address territorial functioning explicitly in the design of user-defined AI agents as a shared resource. For example, designs could either seek to counteract this tendency or lean into it, supporting emerging norms around individual ownership within shared AI environments.
@@ 7.1.2. How many agents? One for fast functionality, many for perspectives and future prompting
Co-writers’ preferences for the number of agents varied, often within the same team, revealing strategic thinking: People weighed the effort of creating multiple agents against the expected benefits, and the nature of those benefits was crucial. Those focused on AI’s functional value (e.g. speed and efficiency) tended to prefer a single agent, while those seeking perspective value (e.g. diverse viewpoints and roles) were more inclined to use multiple agents.
@@
Beyond identifying both values as part of a tradeoff here, our findings reveal how participants related both to prompting: Some linked perspective value to functional benefit by reasoning that role-specific or persona-based agents are a good investment, as they make future prompting more effective. Drawing on prior work, we interpret this as preloading an explicit goal, which reduces ambiguity and may help address prompting challenges (see (Zamfirescu-Pereira et al., 2023)). Concretely, this reflects findings by Schelble et al. (2022) that explicitly shared goals support alignment in human-agent teams, and aligns with Jiang et al. (2023), who observed that clearly assigned roles help users anticipate chatbot behaviour and tailor their input.
In summary, adding to the literature, we show that agent quantity can be more than a technical choice: When left to teams, it becomes part of collaborative deliberation, shaped by anticipated needs around prompting, perspectives, and functionality. Future designs should consider these aspects when deciding whether (or how) to support one vs many agents in a particular collaborative use case.
@@ 7.2.1. Created alone, used together: Emerging norms around shared agents
Taken together, these practices reflect all three functions of common objects in collaborative writing, which Larsen-Ledet et al. (2020) describe as “interim material and outcome of work, a locus of coordination, as well as epistemic and pointing ahead.” We therefore conclude that agent outputs functioned as common objects, in contrast to agent profiles. Since collaborative conventions take time to form (Mark, 2002), it may be faster to establish them around agent output than profiles, as the outputs were embedded in the familiar collaborative material (Bødker et al., 2024) of draft comments. In addition, comments offer clearer actions (e.g. accept, reject, respond) than profiles, which prior work has shown can be difficult to write (Benharrak et al., 2024).
For design, this shows that AI agents can be usefully integrated into established collaboration features such as comments. An interesting design direction is to explore how such integrations can serve as an entry point for teams to establish collaborative conventions around AI use before engaging in the more complex task of creating agents. For example, user-selected comments by a default agent (e.g. liked for their scope, content or tone) could serve as a basis for creating a more specific agent later.
@@ 7.2.2. From agents to tools: Used by the team, not part of it
A key insight from our combined quantitative and qualitative analysis is that participants did not regard the created agents as collaborators.
This was reflected in several ways. First, participants applied authorship and ownership norms to agents and integrated them into structures of responsibility and coordination, thus treating them as material or tools, not as independent. Also in line with this mental model, participants strongly favoured manual control of agent behaviour and welcomed our design decision to surface output as comments to manually accept or reject. Finally, agents were used for familiar tool-like tasks, such as grammar checks, even when other uses were also explored.
These findings reveal that collaborative writing – when adding shared AI agents – remains a fundamentally human, social activity with interpersonal norms and expectations. We thus contribute to the literature on team-AI interaction by demonstrating how shared agents are appropriated through existing socio-technical structures (e.g. territoriality, control, coordination) and that emerging practices around these structures inform users’ mental models of shared AI in collaborative work.
@@ 7.3.1. From comments to contextual conversations
Participants rarely engaged in multi-turn conversations in the comments (Section 6.3.3). They rather used comments as lightweight annotations or as a mechanism for delegating tasks. For extended discussions, participants reported switching to other channels, such as chats and calls.
This behaviour has implications for the design of closely integrated agents in collaborative documents: Rather than serving as a conversational space, comments might be designed as contextualizing connections between selected document parts and conversations. For example, comments could include a “continue in chat” button as a gateway to open a document chat or a dedicated team communication tool such as Slack or Mattermost. Unlike comment threads, these UIs provide more space for elaboration and multi-turn discussions, and better align with users’ existing communication habits. Crucially, these tools would have to integrate the agents and link back to the comments to allow for context-centered, shared AI use.
@@ 7.3.2. Focus- and collaboration-aware agent initiative
These agent-initiated comments were accepted far less often than user-initiated ones (Section 6.3.3), as participants reported feeling cognitively and visually overwhelmed by these document-wide AI interventions. This echoes issues with verbosity and constraints for LLM-generated output in related work (Benharrak et al., 2024; Kabir et al., 2024; Sun et al., 2023), which could be pragmatically addressed with further prompt engineering. While further prompt engineering was beyond the scope of this study, we explored constraining agents more using the existing UI based on participants’ feedback. For example, we found that it is possible to solve the issue of AI suggestions including “opening text” (e.g. “Sure, I’ve drafted this for you”) by adding a note to the profile with “Only output direct draft text, nothing else (e.g. no responses, confirmations, etc.).”
@@
Beyond that, conceptually, we suggest to explore a new middle ground: focus- and collaboration-aware agent initiative. Rather than ignoring the team’s current focus entirely or relying only on explicit user text selections, agents could adapt their initiative to signals of attention, such as cursor or viewport positions, recent edits, or collaborative activity. This information could guide when and how to display suggestions. For example, in-focus suggestions might appear directly as comments or inline lists, while document-wide suggestions could be offered as a pull interaction: initially hidden and displayed only when needed, for example, by stepping through a list of results linked to specific tasks.
A related idea is to simulate agent attention and giving agents an in-document presence indicator (cf. (Prasongpongchai et al., 2025)), similar to the coloured text cursors for users in synchronous text editing.
7.4. Reflections on Methodology
The present study comes with limitations: Our sample consisted mainly of academic writers (faculty and students) and does not represent the broader population of collaborative writers. Nevertheless, it spans a range of experience levels, and academic writing is a key domain for collaborative text work, often used as a context for new writing tools in HCI and beyond (e.g. (Lee et al., 2024; Strobl et al., 2019)).
Our prototype realised agent behaviour and responses via prompting an LLM. While we iterated on the prompts throughout development, we do not claim an “optimal” technical implementation, as our focus in this paper was on deploying the system as a design probe to explore how teams create and interact with such agents.
Although our design probe was available for a week, it was not used every day. Still, norms and perspectives on agent creation and use emerged consistently across groups within this timeframe. A long-term study could examine how collaboration behaviours and conventions evolve over time.
@@ 7.4. Reflections on Methodology
We chose not to include a comparative baseline in favour of maximising participants’ time with the new features in our probe. For a fundamental activity like writing, we assume people can reflect on their usual workflow without reenacting it in a study. Indeed, participants such as $P_{2,1}$, $P_{2,3}$, $P_{3,4}$, $P_{1,5}$, $P_{1,9}$, and $P_{2,12}$ made such comparisons in the interviews (e.g. contrasting our design with their use of ChatGPT). A follow-up study could examine alternative design choices in comparison to our probe and/or other tools (cf. (Mackay and McGrenere, 2025)).
@@ 8. Conclusion
Current AI writing tools are designed for individual use, with little consideration for collaborative settings. To address this gap, we built a functional prototype and examined how co-writers create and interact with shared AI agents during collaborative writing.
We found a clear divide: Agent profiles were treated as personal territory, while created agents and their outputs became shared resources. Agents were used with a preference for manual control rather than autonomous behaviour. This reflects how teams incorporate AI into existing norms of authorship, control, and coordination, rather than treating agents as equal team members. Teams also deliberated over how many agents to create, weighing the efficiency of a single, general-purpose agent against the value of multiple role-specific agents for future prompting and diverse perspectives.
@@ Acknowledgements.
Appendix A User Study
A.1. Participants
@@ A.2. Questionnaire Results
[…]
A three-column table listing the ratings on the CSI factors. Factor Average Factor Count Average Factor Score Collaboration 15.13 2.43 Enjoyment 14.13 2.53 Exploration 13.23 2.43 Expressiveness 12.30 1.97 Immersion 9.40 1.60 Results Worth Effort 13.70 4.03
@@
Appendix B Prompts
B.1. Agent initialization for conversation

### v2

@@ 6.3.1. Agent profiles
Tasks were most frequently concerned with revising a document (38), followed by drafting (20) and outlining (5). Users’ primary intents in task descriptions were efficiency (32), improving the text / readability (22), seeking inspiration (18), and learning (15). The mean text length of task descriptions was 12.46 words (SD=14.04, range: 1-102). Tasks for revising were most often assigned to the general AI agent (14), followed by agents with the roles advisor (10), co-author (8), assistant (6), and editor (5). Tasks for drafting were also most often assigned to the general AI author (14), followed by co-author (3), advisor (2), and editor (2). Tasks for outlining were assigned most often to agents with the role co-author (3) and the general AI author (2).
@@
Tasks in comments most frequently concerned drafting (36) and revising (34), and less frequently outlining (4). Users’ primary intents in comment delegations were inspiration (35), improving the text / readability (24), efficiency (21), and learning (17). The mean text length was 12.23 words (SD=11.75, range: 0-66). Note that we removed the mentioning of agents before computing length (e.g. “@aiRephraser”). Ad-hoc tasks for drafting were assigned most often to agents with the roles co-author (19), advisor (8), the general AI author (8), and editor (2). For revising, ad-hoc tasks were assigned most often to agents with the roles co-author (12), advisor (6), editor (6), the general AI author (5), and agents with an assistant role (5). Ad-hoc tasks for outlining were assigned to agents with an editor role (3), and the general AI author (1).
@@ 7. Discussion
We discuss the findings on our two research questions and further aspects. Fig. 6 provides a high-level overview.
@@ 7.1.1. Agent profiles as territory: Each agent belongs to its creator
As a key insight, although our design allowed anyone to edit any agent profile (Section 4), participants almost never did so (Section 6.1.2). Here, we dissect this in detail:
Overall, this behaviour echoes prior work on collaboration avoidance, such as role dynamics and accountability and credit of contribution described by Wang et al. (2017). Related concerns in our study included the strongly perceived roles of “creator” and “user” for each agent. To examine the reasons for these roles, we refer to what Larsen-Ledet and Korsgaard (2019) describe in their dimensions of writing territoriality as linkage with “place” (incl. local expertise) and interpersonal function. In this light, participants’ behaviour and comments suggest that the creator of an agent is respected, for interpersonal reasons and/or responsibility as its local expert.
The findings advance our understanding of territorial functioning, extending it to shared agents, by contrasting these with shared UI in general: First, user-defined agents result in lasting artefacts (the profiles) that not only show one user’s input to others but also reveal something about the user’s needs and preferences. Thus, agent profiles add a new “place” of text relevant to the interpersonal dimension. People’s rationales for not editing others’ agent profiles show this by referring to familiarity and user representation (Section 6.2.1). In contrast, input via shared tool-like UI is typically ephemeral and thus provides no such “place”, such as when clicking a button. Second, agents add new UI opportunities for future task-related activation by others, in contrast to lasting yet “passive” UI changes related to the document (e.g. formatting settings). In summary, created agents and tasks could be seen as user-defined UI extensions that (1) reveal something about one person’s needs and preferences which (2) defines functionality for others. Adding to the literature on territorial functioning (Larsen-Ledet and Korsgaard, 2019), we thus provide a first account of how this plays out if not only the document writing is shared but also the writing of prompts/profiles for agents: This new type of writing also becomes potential territory, for familiar interpersonal reasons.
A related perspective examines agents as support actors: Gero et al. (2023) identified availability, trust, and individuality as aspects that shape a writer’s perception of a support actor (human or AI). Through this lens, our findings imply that participants transferred these values to the new setting of a shared AI support actor that is always available: Participants included individual familiarity and trust (e.g. “I would not like [others to edit it] if I’m not seeing this”, $P_{2,7}$) in their rationales for “owning” an agent and respecting such ownership of others.
These insights invite future work to address territorial functioning explicitly in the design of user-defined AI agents as a shared resource. For example, designs could either seek to counteract this tendency or lean into it, supporting emerging norms around ownership within shared AI environments.
@@ 7.1.2. How many agents? One for fast functionality, many for perspectives and future prompting
Co-writers’ preferences for the number of agents varied, often within the team, revealing strategic thinking. What drives their preferences is how they weighed the effort of creating multiple agents against the expected benefits. Those focused on AI’s functional value (e.g. speed, efficiency) tended to prefer a single agent, while those seeking perspective value (e.g. diverse viewpoints and roles) were more inclined to use multiple agents.
@@
Beyond identifying both values as part of a tradeoff here, our findings reveal how participants related both to prompting: Some linked perspective value to functional benefit by reasoning that role-specific or persona-based agents are a good investment, as they make future prompting more effective. Drawing on prior work, we interpret this as preloading an explicit goal, which reduces ambiguity and may help address prompting challenges (Zamfirescu-Pereira et al., 2023). Concretely, this reflects findings by Schelble et al. (2022) that explicitly shared goals support alignment in human-agent teams, and aligns with Jiang et al. (2023), who observed that clearly assigned roles help users anticipate chatbot behaviour and tailor input.
In summary, adding to the literature, we show that agent quantity can be more than a technical choice: When left to teams, it becomes part of collaborative deliberation, shaped by anticipated needs around prompting, perspectives, and functionality. Future designs should consider these aspects when deciding whether (or how) to support one vs many agents in a collaborative use case.
@@ 7.2.1. Created alone, used together: Emerging norms around shared agents
Together, these practices reflect all three functions of common objects in collaborative writing, which Larsen-Ledet et al. (2020) describe as “interim material and outcome of work, a locus of coordination, as well as epistemic and pointing ahead.” Hence, agent outputs functioned as common objects, in contrast to agent profiles. Since collaborative conventions take time to form (Mark, 2002), it may be faster to establish them around agent output than profiles, as outputs were embedded in the familiar collaborative material (Bødker et al., 2024) of comments. In addition, comments offer clearer actions (e.g. accept, reject, respond) than profiles, which prior work has shown can be difficult to write (Benharrak et al., 2024).
For design, this shows that AI agents can be usefully integrated into an established collaboration feature (comments). An interesting direction is to explore how such integrations can serve as an entry point for teams to establish collaborative conventions around agents before engaging in the more complex task of creating them. For example, user-selected comments by a default agent (e.g. liked for their scope, content or tone) could serve as a basis for creating a more specific agent later.
@@ 7.2.2. From agents to tools: Used by the team, not part of it
A key insight from our combined quantitative and qualitative analysis is that participants did not regard agents as collaborators: First, participants applied authorship and ownership norms to agents and integrated them into structures of responsibility and coordination, thus treating them as material or tools, not as independent. Moreover, agents were used for familiar tool-like tasks, such as grammar checks, even when other uses were also explored. Finally, although our design enabled participants to give AI initiative and more agency through autonomous triggers, participants favoured manual control of agent behaviour and valued the comment integration with controls to accept or reject.
These findings reveal that collaborative writing – when adding shared agents with the ability to initiate or join collaborative comments – remains a fundamentally human, social activity with interpersonal norms and expectations. We thus contribute to the literature on team-AI interaction by demonstrating how shared agents are appropriated through existing socio-technical structures (e.g. territoriality, control, coordination) and that emerging practices around these structures inform users’ mental models of shared AI in collaborative work.
@@ 7.3.1. From comments to contextual conversations
Participants rarely engaged in multi-turn conversations in the comments (Section 6.1.4), although they valued the agent integration with comments highly (Section 6.2.5). They used comments as lightweight annotations or for delegating tasks. However, they had a need for longer discussions as well: For these, participants reported switching to other channels, such as chats and calls.
This behaviour has implications for the design of closely integrated agents in collaborative documents: Rather than serving as a conversational space, comments might be designed as contextualising connections between selected document parts and conversations. For example, comments could include a “continue in chat” button as a gateway to open a document chat or a dedicated communication tool such as Slack or Mattermost. Unlike comment threads, these UIs provide more space for elaboration and multi-turn discussions, and align with users’ existing communication habits. These tools would have to integrate agents and link back to the comments to allow for context-centered, shared AI use.
@@ 7.3.2. Focus- and collaboration-aware agent initiative
Our results show that agent-initiated comments were accepted far less often than user-initiated ones (Section 6.1.4), as participants reported feeling cognitively and visually overwhelmed by these document-wide AI interventions (Section 6.2.5). This echoes issues with verbosity and constraints for LLM-generated output in related work (Benharrak et al., 2024; Kabir et al., 2024; Sun et al., 2023), which could be pragmatically addressed by refining prompts. While further prompt engineering was beyond our scope here, we explored constraining agents more using the existing UI based on participants’ feedback. For example, we found it possible to solve the issue of AI suggestions including “opening text” (e.g. “Sure, I’ve drafted this for you”) by adding a note to the profile with “Only output direct draft text, nothing else (e.g. no responses, confirmations, etc.).”
Conceptually, such fixes still require a developer-chosen scope for agent comments, or alternatively, require users to specify such a scope per task, which adds to the demands of setup and prompting (cf. Section 6.2.5 and (Benharrak et al., 2024; Zamfirescu-Pereira et al., 2023)). As a potential alternative to shift that decision-making to the system, we suggest exploring a new middle ground: focus- and collaboration-aware agent initiative. This also fits to participants rationales for preferring manual controls, which revealed the desire for keeping AI responses in view, to not miss both (1) expected results (“It works, but I have seen no results”, $P_{2,5}$) and (2) unexpected ones (“You’re always afraid that some changes will happen to your documents”, $P_{1,7}$).
@@
Concretely, rather than ignoring the team’s current focus or relying on explicit user text selections, agents could adapt their initiative to signals of attention, such as cursor or viewport positions, recent edits, or collaborative activity. This could guide when and how to display suggestions. In-focus suggestions might appear as comments or inline lists, while document-wide suggestions could be offered as a pull interaction: initially hidden and displayed only when needed, for example, by stepping through a list of results linked to specific tasks. A related idea is to simulate agent attention and giving them an in-document presence indicator (cf. (Prasongpongchai et al., 2025)), similar to the coloured text cursors for users.
7.4. Limitations and Ideas for Future Work
@@
This study comes with the typical limitations of a qualitative exploration. For example, we chose not to include a comparative baseline in favour of maximising participants’ time with the new probe features. For a fundamental activity like writing, we assume people can reflect on their usual workflow without reenacting it in a study. Indeed, participants made such explicit comparisons (Section 6.2.5). Future work could examine alternative design choices in comparison to our probe and/or other tools (cf. (Mackay and McGrenere, 2025)). For example, the integration of agent responses into collaborative comments could be compared to a baseline with a typical chatbot in a sidebar. Moreover, a quantitative follow-up study could compare a design that supports agent ownership explicitly (e.g. “Anna’s agent” shown next to the agent’s name if Anna created it) against a design that tries to counteract these identified ownership perceptions (e.g. requiring input/confirmation from all users before a new agent can be used).
While our sample of academic writers does not represent the broader population, it spans a range of experience levels, and academic writing is a key domain for collaborative text work, often used as a context for new tools (e.g. (Lee et al., 2024; Strobl et al., 2019)). While different relationships, including hierarchies, might impact teamwork, our sample did not surface any patterns in this regard. Note that except for group 7, people in each team knew each other beforehand. Future work could study a larger sample to systematically compare use of shared agents by teams with different characteristics in this regard.
While we iterated on the prompts in our prototype throughout development, we do not claim an “optimal” implementation, as our focus was on deploying the system as a technology probe (Feng et al., 2025; Hutchinson et al., 2003). A probe deployment is limited. Still, norms and perspectives on agent creation and use emerged consistently across groups within a week. A long-term study could examine evolving collaboration over time.
@@ 8. Conclusion
Current AI writing tools are designed for individual use, with little consideration for collaborative settings. To address this gap, we built a functional prototype and examined how co-writers create and interact with shared AI agents during collaborative writing. Our qualitative study sheds light on values and demands of three new design features (Section 6.2.5): agent profiles, agent tasks, and integration of agent responses into comments, via both user- and system-initiated triggers.
Overall, we found a clear divide to be considered in future research and design: Agent profiles were treated as personal territory, while created agents and their outputs became shared resources. Agents were used with a preference for manual control rather than autonomous behaviour. This reflects how teams incorporate AI into existing norms of authorship, control, and coordination, rather than treating agents as equal team members, although our design also would have supported choosing greater agency for AI if participants had desired it. Teams also deliberated over how many agents to create, weighing perceived efficiency for a single, general-purpose agent against the value of multiple role-specific agents for future prompting and diverse perspectives.
@@ Acknowledgements.
Appendix A Formative Survey to Inform Autonomous Triggers
We conducted a short, formative survey to gather ideas for autonomous triggers that writers expect to be useful and relevant. We recruited 16 participants through our academic network (7 women, 9 men, age 22 - 38 years), covering students and researchers from technical and social science backgrounds. The questionnaire introduced the idea of agents acting in a shared text document and then asked three open questions:
  • What features would make an AI agent useful for you when collaborating on a document with other people?
  • Which autonomous tasks would you ask an agent to do for you?
  • Define rules that specify when the agent should autonomously do its job to suggest or make changes.
The mentioned tasks that participants expected to be suitable for delegation to AI agents included grammar and spelling correction, translation, text drafting, idea generation, content expansion, citation management, information retrieval, text formatting, and document processing according to custom criteria.
Moreover, participants proposed a diverse range of triggers. We grouped them into three categories:
  • Time-based, such as: “every 30 minutes”, “in 1 minute after I stop typing”, “whenever I haven’t typed anything within the past 30 seconds”
  • User action-based, such as: “when I select text or delete paragraph”, “when I finish a sentence or paragraph or page”, “whenever I hover over a word”, “when inserting a table or a picture”, “when the user reopens the document after a long time of inactivity”, “when I’m working with multiple people”.
  • Event or pattern-based, such as: “when a section has been completed”, “whenever you find a critical error like plagiarism”, “after completing the first draft”, “whenever my text is too repetitive”.
Several of the proposed rules express hesitation regarding autonomous AI behaviour, such as “Never start autonomously and use track-change and protocol suggestions to the text.” Some participants preferred AI to act only when explicitly requested to do so; or they did not wish to get suggestions while typing or at early writing stages. Others were open to the idea of assistance in the background, as long as agents did not disrupt them or update text without approval. These concerns fit to the later findings in the user study, which revealed an overall preference for manually triggering agents (Section 6.2.5).
The suggested rules informed our set of triggers in the prototype (Section 4.4). Concretely, we decided to offer simple time-based triggers (short intervals, inactivity) and event-based triggers (document saved, collaborative edits, users offline) that do not require further parameter choices from users. Note that “in short intervals” is very flexible: It can be used to setup tasks that realise background assistance reacting to text “all the time” (e.g. checking grammar). Thus, it also covers the semantic or stylistic rules suggested by participants (e.g. “Suggest alternative sentences when I stray too far from formal style”). In this way, the implemented triggers cover 20 of the 29 rules proposed by participants. Note that the remaining nine rules described manual triggers or a preference for no automated triggering. This is covered by our prototype’s support for manually triggering AI (e.g. selecting text to run a specific prompt on that part).
Appendix B User Study
B.1. Participants
@@ B.2. Questionnaire Results
A three-column table listing the ratings on the CSI factors.
Factor Average Factor Count Average Factor Score
Collaboration 15.13 2.43
Enjoyment 14.13 2.53
Exploration 13.23 2.43
Expressiveness 12.30 1.97
Immersion 9.40 1.60
Results Worth Effort 13.70 4.03
@@
B.3. Further logging results
Table 6. Count of interaction switches per group, plus percentages of switches to humans vs AI agents as an actor (i.e. events when AI adds a comment, and AI agent involvement via suggestion acceptance events). Note that one switch does not equal one action because e.g. a user might make a sequence of many edits before another user (or agent) acts. Similarly, an agent might add more than one comment in sequence, e.g. as a result of running a document task. Overall, these values contextualise our findings and show that teams in our sample had varying collaboration styles in terms of how much their work on the shared document interleaved with each other and with AI agents.
A four-column table listing the interaction switches of each group.
Group Interaction switches
Total to human (%) to AI (%)
1 580 96.4 3.6
2 99 94.9 5.1
3 35 14.3 85.7
4 - - -
5 42 76.2 23.8
6 220 78.2 21.8
7 136 45.6 54.4
8 10 40.0 60.0
9 337 83.1 16.9
10 20 20.0 80.0
11 18 11.1 88.9
12 69 88.4 11.6
13 292 81.2 18.8
14 990 99.5 0.5
Appendix C Prompts
This appendix lists the prompts as used in our implementation for the study, at the end of our development process of two years. They were iterated on throughout discussions and informal testing in our research group and with colleagues with backgrounds in HCI and NLP. In general, we started with simple prompts that include the main instruction and added further information such as the “guidelines” based on observed shortcomings in our testing.
C.1. Agent initialization for conversation